# Friends not Foes: Strong Correlation between Inner Super-Earths and Outer Gas Giants

### v1

@@ Abstract
The connection between outer gas giants and inner super-Earths reflects their formation and evolutionary histories. Past work exploring this link has suggested a tentative positive correlation between these two populations, but these studies have been limited by small sample sizes and in some cases sample biases. Here we take a new look at this connection with a sample of 184 super-Earth systems with publicly available radial velocity data and fully resolved outer gas giants. We calculate the frequency of outer gas giants (GG) in super-Earth (SE) systems, dividing our sample into metal-rich ([Fe/H] 0) and metal-poor ([Fe/H]0) hosts. We find P(GGSE, [Fe/H]0) = 28.0 and P(GGSE, [Fe/H]0) = 4.5. Comparing these conditional occurrence rates to field giant occurrence rates from Wittenmyer et al. (2020), we show that there is a distinct positive correlation between inner super-Earths and outer gas giants for metal-rich host stars at the 2.5 level, but this correlation disappears for metal-poor systems. We further find that, around metal-rich stars, the GG/SE correlation enhances for systems with giants that are more distant (beyond 3 AU) and/or more eccentric (), while gas giant multiplicity does not appear to affect the level of correlation. Such trends again disappear around metal-poor stars with the exception of systems of multiple giants in which we observe an anti-correlation. Our findings highlight the critical role metallicity (disk solid budget) plays in shaping the overall planetary architecture.
@@ 1 Introduction
The connection between inner super-Earths and outer gas giants was first explored in Zhu & Wu (2018a) and Bryan et al. (2019). Both studies found that outer gas giants do not hinder the formation of super-Earths, and instead show a positive correlation between the two populations. In addition, most of the gas giant companions were found around metal-rich host stars, indicating metallicity is an important factor driving this correlation. Subsequent theoretical work explains this excess of cold Jupiters in super-Earth systems by showing that for systems with enough solids to produce a gas giant, they naturally deliver enough material to the inner system for super-Earths to grow (e.g., Chachan et al., 2022; Chachan & Lee, 2023) since most of the solids are lost to radial drift rather than being coagulated into planetary cores at large orbital distances (e.g., Lin et al., 2018).
While there have been follow-up efforts that also find a positive correlation between gas giants and small planets (e.g. Herman et al., 2019; Rosenthal et al., 2022), some have reported a neutral or even negative correlation (e.g. Bonomo et al., 2023). A resolution to this debate was presented in Zhu (2023), arguing that interpreting this correlation requires considering the host star metallicity distribution of the sample of systems. Namely, since the vast majority of the gas giants in all samples considered to date are found around metal-rich stars, having a sample of systems dominated by metal-poor stars (as was the case in Bonomo et al. (2023)) can bias the strength of the correlation. Zhu (2023) showed that when only considering the metal-rich systems in Zhu & Wu (2018a), Bryan et al. (2019), and Bonomo et al. (2023), all three find consistent positive correlations between super-Earths and cold gas giants.
@@
In this paper, we explore the connection between inner super-Earths and outer gas giants with a sample of 184 systems, at least three times larger than previous samples considered, with the orbit of giant(s) resolved by radial velocity data. In Section 2, we describe how these systems were selected. Section 3 details our computation of sensitivity maps and completeness correction for each system, and Section 4 shows how we incorporate those maps into our occurrence rate calculations to compute the frequency of cold gas giants in super-Earth systems that are metal rich P(GGSE, [Fe/H]0) and metal poor P(GGSE, [Fe/H]0). To assess the presence and strength of a correlation, we compare our conditional probabilities to gas giant occurrence rates P(GG[Fe/H]0) and P(GG[Fe/H]0) from Wittenmyer et al. (2020) (hereafter denoted W20) in Section 5. Section 6 explores how gas giant properties shape this correlation, and Section 7 summarizes our findings, placing them in the context of formation theory.
@@ 2 Observations
We select systems with the following five criteria: the system has 1) publicly available radial velocity (RV) datasets, 2) at least one confirmed super-Earth planet (1-20 M; 1-4 R), 3) host star mass greater than 0.6 M to exclude M-dwarfs, 4) RV datasets with baselines greater than one year and more than 20 data points, and 5) an RV semi-amplitude K1m/s in the case of systems with super-Earths discovered using RVs to minimize the inclusion of false positive planets. We exclude M-dwarfs from this analysis for several reasons. Given the selection criteria above, there are 78 systems with host star masses 0.6 M that we could consider. However, out of the 203 stars in W20, only 8 are M-dwarfs. We thus leave the assessment of the gas giant/super-Earth correlation around M-dwarfs to subsequent work with a more appropriate comparison sample. We additionally exclude a handful of systems that do not have published metallicity measurements. This yields a total of 184 systems (see Table A.1). 109 of these systems have one or more super-Earth initially discovered via the transit method (hereafter labeled ‘transit sample’), and 75 systems have one or more super-Earth initially discovered with RVs (labeled ‘RV sample’). As described in the following section, we initially consider these samples separately due to the significant average difference in sensitivities to distant gas giants.
Of the systems in the transit sample, 56 are metal-rich [Fe/H]0, and 53 are metal-poor [Fe/H] 0. Similarly, the RV sample has 39 metal-rich systems and 36 metal-poor ones. We consider a system to host a gas giant (0.520 M) if there is a confirmed, resolved gas giant in the system. Given this criteria, there are 19 systems in the transit sample hosting at least one gas giant, 17 of which are metal-rich. The RV sample has 11 systems with one or more gas giants, 10 of which are metal rich. Properties of the gas giants in this sample are shown in Table A.2. We show host star properties of our sample in Figure 1.
@@ 3 Sample Sensitivity Maps
With a heterogeneous sample such as that considered here, individual system sensitivities to outer gas giants can vary widely due to differences in time baseline, observing cadence, number of data points, and measurement precision. For each system, we quantify sensitivity by calculating completeness as a function of mass and semi-major axis given the system’s RV dataset (see Table A.1 for references).111In the case where there is more than one public RV dataset that has a time baseline 1 year and more than 20 data points, we select the dataset that has higher completeness. We start with a 5050 grid in mass and semi-major axis evenly spaced in log space from 0.330 M and 0.330 AU. We inject 50 simulated planets into each grid box, each time drawing a mass and semi-major axis value from a uniform distribution across each grid box, an eccentricity from the distribution, an inclination from a uniform distribution in , and remaining orbital elements from uniform distributions.
@@
Figure A.1 in the appendix presents average completeness maps for the RV super-Earth sample (top) and the transit super-Earth sample (bottom). Detected gas giants in each sample are overplotted. The reason why we initially separate these samples in our analysis is evident: on top of variations amongst individual systems, on average the RV sample systems have significantly higher sensitivities, driven by as much as an order of magnitude more data points and longer time baselines in comparison to the transit sample. Despite these differences, confirmed gas giants are in regions of parameter space where both transit and RV samples are on average close to, if not at, 100 completeness. This motivates our mass and semi-major axis ranges considered in our occurrence rate calculations in the following sections, namely 0.520 M and 110 AU. Individual sensitivity maps allow us to account for completeness to gas giants in this range on a system-by-system basis in these calculations, and confirm that these sensitivity differences do not bias our results.
@@ 4 Consistent Occurrence Rates P(GGSE) for Transit and RV Samples
To explore the connection between super-Earths and Jupiters, we start by calculating the frequency of gas giants (GG) in systems with inner super-Earths (SE), P(GGSE). We define these gas giants as planets with mass 0.520 M and orbital distance 110 AU. Previous work has demonstrated the importance of host star metallicity to this GG/SE connection (Zhu, 2023). We split our sample into metal-rich systems [Fe/H]0, and metal-poor systems [Fe/H]0 to verify the sensitivity of this correlation to metallicity.
We determine P(GGSE, [Fe/H]0) and P(GGSE, [Fe/H]0) using a general binomial distribution. If all of our systems had 100 completeness across our entire integration range (0.520 M and 110 AU), the posterior distribution given number of detections and total number of systems is:
where is the beta function, = +1, and = -+1. To correct for individual system sensitivities to gas giants, we used the completeness maps we computed for each system. We sum over the 0.520 M and 110 AU range and determine the average completeness for the set of systems considered. The coefficient then becomes = , where is the total number of systems modified by the sample completeness.
We initially consider the RV and transit samples separately to test whether the average differences in system sensitivity impact the resulting occurrence rates. There are 56 metal-rich transit sample systems, 14 of which host one or more gas giants 0.520 M and 110 AU, 53 metal-poor transit sample systems with 2 hosting gas giants, 39 metal-rich RV sample systems with 10 hosting gas giants, and 36 metal-poor RV sample systems with just one hosting a gas giant. We find that the frequencies of gas giants in super-Earth systems for both metal poor and metal rich systems are consistent when comparing the RV and transit samples: P(GGSE, [Fe/H])0) = 28.8 (+6.5 -6.0) for the transiting sample and P(GGSE, [Fe/H]0) = 27.8 (+7.5 -6.8) for the RV sample, consistent to . Similarly, the metal poor RV and transit samples are consistent at the level, with P(GGSE, [Fe/H])0) = 5.5 (+3.9 -2.7) for the transit sample and P(GGSE, [Fe/H]0) = 4.8 (+4.4 -3.8) for the RV sample. We can therefore move forward with a combined RVtransit sample occurrence rate for the rest of the analyses. For this combined sample, we calculate conditional occurrence rate P(GGSE, [Fe/H]0) = 28.0 (+4.9 -4.6), and a substantially lower frequency for metal poor systems, P(GSE, [Fe/H]0) = 4.5 (+2.6 -1.9) (see Table 1).
@@ 5 A Strong Positive Correlation Between SE/GG For Metal Rich Systems
To assess whether there is a positive, negative, or no correlation between super-Earths and gas giants in our sample, we need to compare our conditional occurrence rate P(GGSE) to the field occurrence of gas giants P(GG). Here we use the sample of systems from Wittenmyer et al. (2020) (W20). This work presents the frequency of gas giants P(GG) using RV data from the Anglo-Australian Planet Search, leveraging the 18-year archive of data for 203 stars. In this paper we use all systems in the W20 sample except ones that are M-dwarfs and ones that do not have published stellar metallicities. The W20 sample considered here has 107 metal-rich stars, 14 of which host gas giants (0.520 M and 110 AU); and 87 metal-poor stars, 5 of which have one or more gas giants (same definition as the metal-rich sample). Note again that we exclude M-dwarfs from this analysis to ensure consistent comparison. We leave the assessment of the GG/SE correlation around M-dwarfs to subsequent work.
@@
To compare our value of P(GGSE) to the W20 value for P(GG), it is critical to 1) use the same definition of a gas giant, namely mass and semi-major axis ranges considered, and 2) use comparable estimates for sensitivity corrections. In short, W20 carries out injection/recovery simulations where they assess whether a planet has been “detected" based on whether a Lomb-Scargle periodogram can recover that signal with a false alarm probability (FAP) 1 (see W20 for details). To test whether our method presented in Section 3 is comparable, we recalculate sensitivity maps for all of our systems using the W20 LS method. We find P(GGSE, [Fe/H]0) = 26.3 (+4.6 -4.4) and P(GGSE, [Fe/H]0) = 4.2 (+2.4 -1.8), consistent with our original values to 0.3 and 0.1 respectively. We can thus move forward with our original completeness estimates for P(GGSE).
@@
With consistent sensitivity corrections, we calculate P(GG[Fe/H]0) and P(GG[Fe/H]0) for the W20 sample, using consistent definition of a gas giant (0.520 M and 110 AU). Figure 2 and Table 1 show our results. For metal-rich systems, we find a distinct positive correlation between gas giants and super-Earths at the 2.5 level. There are more gas giants in metal-rich super-Earth systems than you would expect just based on chance. In contrast, for metal poor systems the positive correlation disappears, and may in fact go tentatively negative. These results highlight the importance of this metallicity dimension when interpreting the SE/GG correlation. If a sample is biased towards low metallicities or there is a mixture of low and high, the strength of the correlation can be artificially underestimated.
Test | BL [Fe/H]0 | BL [Fe/H]0 | W20 [Fe/H]0 | W20 [Fe/H]0 | [Fe/H]0 | [Fe/H]0 |
(1) Default | 28.0 (+4.9 -4.6) | 4.5 (+2.6 -1.9) | 13.5 (+3.5 -3.0) | 6.4 (+2.9 -2.3) | 2.5 | -0.5 |
(2) Close GG | 21.4 (+4.4 -4.0) | 3.1 (+2.2 -1.5) | 12.6 (+3.4 -2.9) | 6.4 (+2.9 -2.3) | 1.7 | -0.8 |
(3) Distant GG | 18.3 (+4.4 -3.9) | 2.1 (+2.0 -1.2) | 5.2 (+2.4 -1.9) | 3.0 (+2.2 -1.5) | 3.0 | -0.4 |
(4) Dynamically Hot | 19.2 (+4.3 -3.9) | 2.0 (+1.9 -1.0) | 7.1 (+2.7 -2.2) | 4.2 (+2.4 -1.8) | 2.6 | -0.8 |
(5) Dynamically Cold | 11.6 (+3.6 -3.0) | 3.1 (+2.2 -1.5) | 8.0 (+2.8 -2.3) | 4.2 (+2.4 -1.8) | 0.9 | -0.4 |
(6) Single GG | 17.3 (+4.0 -3.6) | 4.1 (+2.4 -1.7) | 11.7 (+3.3 -2.8) | 4.2 (+2.4 -1.8) | 1.7 | 0.0 |
(7) Multi GG | 8.0 (+3.0 -2.5) | 2.0 | 2.5 (+1.8 -1.2) | 3.0 (+2.2 -1.5) | 1.8 | -1.2 |
@@
Note. — (1) Default: GG 1–10AU, 0.5–20M; (2) Close GG: 0.1–3AU, 0.5–20M; (3) Distant GG: 3–10AU, 0.5–20M; (4) Dynamically Hot: default GG with eccentricities 0.2; (5) Dynamically Cold: default GG with eccentricities 0.2; (6) Single GG: systems with a single gas giant in the default GG range; (7) Multi GG: systems with multiple gas giants in the default GG range.

### v2

@@ Abstract
The connection between outer gas giants and inner super-Earths reflects their formation and evolutionary histories. Past work exploring this link has suggested a tentative positive correlation between these two populations, but these studies have been limited by small sample sizes and in some cases sample biases. Here we take a new look at this connection with a sample of 184 super-Earth systems with publicly available radial velocity data and resolved outer gas giants. We calculate the frequency of outer gas giants (GG) in super-Earth (SE) systems, dividing our sample into metal-rich ([Fe/H] 0) and metal-poor ([Fe/H]0) hosts. We find P(GGSE, [Fe/H]0) = 28.0 and P(GGSE, [Fe/H]0) = 4.5. Comparing these conditional occurrence rates to field giant occurrence rates from Rosenthal et al. (2021), we show that there is a distinct positive correlation between inner super-Earths and outer gas giants for metal-rich host stars at the 2.7 level, but this correlation disappears for metal-poor systems. We further find that, around metal-rich stars, the GG/SE correlation enhances slightly for systems with giants that are more distant (beyond 3 AU), more eccentric (), and/or in multi-gas giant systems. Such trends disappear around metal-poor stars with the exception of systems of multiple giants in which we observe a tentative anti-correlation. Our findings highlight the critical role metallicity (disk solid budget) plays in shaping the overall planetary architecture.
@@ 1 Introduction
The connection between inner super-Earths and outer gas giants was first explored in Zhu & Wu (2018) and Bryan et al. (2019). Both studies found that outer gas giants do not hinder the formation of super-Earths, and instead show a positive correlation between the two populations. In addition, most of the gas giant companions were found around metal-rich host stars, indicating metallicity is an important factor driving this correlation. Subsequent theoretical work explains this excess of cold Jupiters in super-Earth systems by showing that for systems with enough solids to produce a gas giant, they naturally deliver enough material to the inner system for super-Earths to grow (e.g., Chachan et al., 2022; Chachan & Lee, 2023) since most of the solids are lost to radial drift rather than being coagulated into planetary cores at large orbital distances (e.g., Lin et al., 2018).
While there have been follow-up efforts that also find a positive correlation between gas giants and small planets (e.g. Herman et al., 2019; Rosenthal et al., 2022), some have reported a neutral or even negative correlation (e.g. Bonomo et al., 2023). A resolution to this debate was presented in Zhu (2023), arguing that interpreting this correlation requires considering the host star metallicity distribution of the sample of systems. Namely, since the vast majority of the gas giants in all samples considered to date are found around metal-rich stars, having a sample of systems dominated by metal-poor stars (as was the case in Bonomo et al. (2023)) can bias the strength of the correlation. Zhu (2023) showed that when only considering the metal-rich systems in Zhu & Wu (2018), Bryan et al. (2019), and Bonomo et al. (2023), all three find consistent positive correlations between super-Earths and cold gas giants.
@@
In this paper, we explore the connection between inner super-Earths and outer gas giants with a sample of 184 systems, at least three times larger than previous samples considered, with the orbit of giant(s) resolved by radial velocity data. In Section 2, we describe how these systems were selected. Section 3 details our computation of sensitivity maps and completeness correction for each system, and Section 4 shows how we incorporate those maps into our occurrence rate calculations to compute the frequency of cold gas giants in super-Earth systems that are metal rich P(GGSE, [Fe/H]0) and metal poor P(GGSE, [Fe/H]0). To assess the presence and strength of a correlation, we compare our conditional probabilities to gas giant occurrence rates P(GG[Fe/H]0) and P(GG[Fe/H]0) from Wittenmyer et al. (2020) (hereafter denoted W20) and from Rosenthal et al. (2021) (hereafter denoted R21) in Section 5. Section 6 explores how gas giant properties shape this correlation, and Section 7 summarizes our findings, placing them in the context of formation theory.
@@ 2 Observations
We select systems with the following five criteria: the system has 1) publicly available radial velocity (RV) datasets, 2) at least one confirmed super-Earth planet (1-20 M⊕; 1-4 R⊕), 3) host star mass greater than 0.6 M☉ to exclude M-dwarfs, 4) RV datasets with baselines greater than one year and more than 20 data points, and 5) an RV semi-amplitude K1m/s in the case of systems with super-Earths discovered using RVs to minimize the inclusion of false positive planets. We exclude M-dwarfs from this analysis for several reasons. Given the selection criteria above, there are 78 systems with host star masses 0.6 M⊙ that we could consider. However, out of the 203 stars in W20, only 8 are M-dwarfs. While the fraction of the R21 sample that are M stars is higher (121 M-dwarfs out of 719 stars), given the relevance of the W20 sample to this study we leave the assessment of the gas giant/super-Earth correlation around M-dwarfs to subsequent work. We additionally exclude a handful of systems that do not have published metallicity measurements. This yields a total of 184 systems (see Table A.1). 109 of these systems have one or more super-Earth initially discovered via the transit method (hereafter labeled ‘transit sample’), and 75 systems have one or more super-Earth initially discovered with RVs (labeled ‘RV sample’). As described in the following section, we initially consider these samples separately due to the significant average difference in sensitivities to distant gas giants.
Of the systems in the transit sample, 56 are metal-rich [Fe/H]0, and 53 are metal-poor [Fe/H] 0. Similarly, the RV sample has 39 metal-rich systems and 36 metal-poor ones. We consider a system to host a gas giant (0.520 MJup) if there is a confirmed gas giant (typically with 1 measured RV orbit) in the system. Given this criteria, there are 19 systems in the transit sample hosting at least one gas giant, 17 of which are metal-rich. The RV sample has 11 systems with one or more gas giants, 10 of which are metal rich. Properties of the gas giants in this sample are shown in Table A.2. We show host star properties of our sample in Figure 1.
@@ 3 Sample Sensitivity Maps
With a heterogeneous sample such as that considered here, individual system sensitivities to outer gas giants can vary widely due to differences in time baseline, observing cadence, number of data points, and measurement precision. For each system, we quantify sensitivity by calculating completeness as a function of mass and semi-major axis given the system’s RV dataset (see Table A.1 for references).111In the case where there is more than one public RV dataset that has a time baseline 1 year and more than 20 data points, we select the dataset that has higher completeness. We start with a 5050 grid in mass and semi-major axis evenly spaced in log space from 0.330 MJup and 0.330 AU. We inject 50 simulated planets into each grid box, each time drawing a mass and semi-major axis value from a uniform distribution across each grid box, an eccentricity from the distribution, an inclination from a uniform distribution in , and remaining orbital elements from uniform distributions.
@@
Figure A.1 in the appendix presents average completeness maps for the RV super-Earth sample (top) and the transit super-Earth sample (bottom). Detected gas giants in each sample are overplotted. The reason why we initially separate these samples in our analysis is evident: on top of variations amongst individual systems, on average the RV sample systems have significantly higher sensitivities, driven by as much as an order of magnitude more data points and longer time baselines in comparison to the transit sample. Despite these differences, confirmed gas giants are in regions of parameter space where both transit and RV samples are on average close to, if not at, 100 completeness. This motivates our mass and semi-major axis ranges considered in our occurrence rate calculations in the following sections, namely 0.520 MJup and 110 AU. Individual sensitivity maps allow us to account for completeness to gas giants in this range on a system-by-system basis in these calculations, and confirm that these sensitivity differences do not bias our results.
@@ 4 Consistent Occurrence Rates P(GGSE) for Transit and RV Samples
To explore the connection between super-Earths and Jupiters, we start by calculating the frequency of gas giants (GG) in systems with inner super-Earths (SE), P(GGSE). We define these gas giants as planets with mass 0.520 MJup and orbital distance 110 AU. Previous work has demonstrated the importance of host star metallicity to this GG/SE connection (Zhu, 2023). We split our sample into metal-rich systems [Fe/H]0, and metal-poor systems [Fe/H]0 to verify the sensitivity of this correlation to metallicity.
We determine P(GGSE, [Fe/H]0) and P(GGSE, [Fe/H]0) using a general binomial distribution. If all of our systems had 100 completeness across our entire integration range (0.520 MJup and 110 AU), the posterior distribution given number of detections and total number of systems is:
where is the beta function, = +1, and = -+1. To correct for individual system sensitivities to gas giants, we used the completeness maps we computed for each system. We sum over the 0.520 MJup and 110 AU range and determine the average completeness for the set of systems considered. The coefficient then becomes = , where is the total number of systems modified by the sample completeness.
We initially consider the RV and transit samples separately to test whether the average differences in system sensitivity impact the resulting occurrence rates. There are 56 metal-rich transit sample systems, 14 of which host one or more gas giants 0.520 MJup and 110 AU, 53 metal-poor transit sample systems with 2 hosting gas giants, 39 metal-rich RV sample systems with 10 hosting gas giants, and 36 metal-poor RV sample systems with just one hosting a gas giant. We find that the frequencies of gas giants in super-Earth systems for both metal poor and metal rich systems are consistent when comparing the RV and transit samples: P(GGSE, [Fe/H])0) = 28.8 (+6.5 -6.0) for the transiting sample and P(GGSE, [Fe/H]0) = 27.8 (+7.5 -6.8) for the RV sample, consistent to . Similarly, the metal poor RV and transit samples are consistent at the level, with P(GGSE, [Fe/H])0) = 5.5 (+3.9 -2.7) for the transit sample and P(GGSE, [Fe/H]0) = 4.8 (+4.4 -3.8) for the RV sample. We can therefore move forward with a combined RVtransit sample occurrence rate for the rest of the analyses. For this combined sample, we calculate conditional occurrence rate P(GGSE, [Fe/H]0) = 28.0 (+4.9 -4.6), and a substantially lower frequency for metal poor systems, P(GSE, [Fe/H]0) = 4.5 (+2.6 -1.9) (see Table 1).
@@ 5 A Strong Positive Correlation Between SE/GG For Metal Rich Systems
To assess whether there is a positive, negative, or no correlation between super-Earths and gas giants in our sample, we need to compare our conditional occurrence rate P(GGSE) to the field occurrence of gas giants P(GG). For P(GG), we use the sample of systems from Rosenthal et al. (2021) (R21) and Wittenmyer et al. (2020) (W20). R21 presents the California Legacy Survey, a catalog of radial velocities for 719 stars observed over the course of 30 years. In this paper we take all systems in the R21 sample except M-star hosts (0.6 M⊙) and remaining systems with datasets 20 RV measurements (4 systems). This yields 351 metal rich systems, 43 of which host gas giants (0.520 MJup and 110 AU); and 243 metal-poor systems, 9 of which host gas giants (same definition as the metal-rich sample). Note again that we exclude M-dwarfs from this analysis to ensure consistent comparison. We leave the assessment of the GG/SE correlation around M-dwarfs to subsequent work. In this section we also consider the W20 sample, which comes from the Anglo-Australian Planet Search that leverages the 18-year archive of data for 203 stars. Here we use all systems in the W20 sample except ones that are M-dwarfs and ones that do not have published stellar metallicities. The W20 sample considered here has 107 metal-rich stars, 14 of which host gas giants; and 87 metal-poor stars, 5 of which have one or more gas giants. Despite the smaller sample size, we initially include the W20 P(GG) value in our analysis since previous work examining the inner-outer planet correlation (e.g., Zhu, 2023; Bonomo et al., 2023) used the W20 sample for P(GG) when interpreting their values for P(SE|GG).
@@
To compare our value of P(GGSE) to the R21 and W20 values for P(GG), it is critical to 1) use the same definition of a gas giant, namely mass and semi-major axis ranges considered, and 2) use comparable estimates for sensitivity corrections. For the R21 sample, we calculate individual system sensitivity maps for the whole sample using our methods outlined in Section 3 using the public radial velocities. Since RVs are not public for the W20 sample, we instead compare our sensitivity estimate methods. In short, W20 carries out injection/recovery simulations where they assess whether a planet has been “detected" based on whether a Lomb-Scargle periodogram can recover that signal with a false alarm probability (FAP) 1 (see W20 for details). To test whether our method presented in Section 3 is comparable, we recalculate sensitivity maps for all of our systems using the W20 LS method. We find P(GGSE, [Fe/H]0) = 26.3 (+4.6 -4.4) and P(GGSE, [Fe/H]0) = 4.2 (+2.4 -1.8), consistent with our original values to 0.3 and 0.1 respectively. We can thus move forward with our comparisons of P(GG) and P(GG|SE).
@@
With consistent sensitivity corrections, we calculate P(GG[Fe/H]0) and P(GG[Fe/H]0) for the R21 and W20 samples, using consistent definition of a gas giant (0.520 MJup and 110 AU). Figure 2 and Table 1 show our results. For both metal rich and metal poor systems, the P(GG) frequencies from R21 and W20 are consistent to 1. From here we solely consider the R21 P(GG) occurrence rates throughout the rest of the paper due to the greater statistical leverage stemming from a significantly larger sample size in comparison to W20. For metal-rich systems, we find a distinct positive correlation between gas giants and super-Earths at the 2.7 level. There are more gas giants in metal-rich super-Earth systems than you would expect just based on chance. In contrast, for metal poor systems the positive correlation disappears. These results highlight the importance of this metallicity dimension when interpreting the SE/GG correlation. If a sample is biased towards low metallicities or there is a mixture of low and high, the strength of the correlation can be artificially underestimated.
Given the size of our sample and that of R21, we break both samples up into three metallicity bins to explore the dependence of the correlation on metallicity further: 1) [Fe/H]-0.1; 2) -0.1[Fe/H]0.1; 3) [Fe/H]0.1. Figure 3 shows that while there is no significant correlation between super-Earths and GG around stars in the first two bins, the most metal-rich bin has an even stronger positive correlation between the two populations, at the 3.4 level.
Test | BL [Fe/H]0 | BL [Fe/H]0 | R21 [Fe/H]0 | R21 [Fe/H]0 | [Fe/H]0 | [Fe/H]0 |
(1) Default | 28.0 (+4.9 -4.6) | 4.5 (+2.6 -1.9) | 14.3 (+2.0 -1.8) | 5.0 (+1.6 -1.3) | 2.7 | -0.2 |
(2) Close GG | 19.2 (+4.3 -3.9) | 3.1 (+2.2 -1.5) | 9.5 (+1.7 -1.5) | 3.8 (+1.4 -1.1) | 2.3 | -0.3 |
(3) Distant GG | 18.3 (+4.4 -3.9) | 2.1 (+2.0 -1.2) | 7.4 (+1.5 -1.3) | 2.5 (+1.2 -0.9) | 2.7 | -0.2 |
(4) Dynamically Hot | 20.1 (+4.5 -4.0) | 2.0 (+1.9 -1.2) | 7.1 (+1.5 -1.3) | 1.6 (+1.0 -0.7) | 3.0 | 0.3 |
(5) Dynamically Cold | 12.1 (+3.7 -3.2) | 3.2 (+2.3 -1.6) | 7.4 (+1.5 -1.3) | 2.9 (+1.2 -1.0) | 1.3 | 0.2 |
(6) Single GG | 17.3 (+4.0 -3.6) | 4.1 (+2.4 -1.7) | 10.4 (+1.7 -1.5) | 3.1 (+1.2 -1.0) | 1.7 | 0.5 |
(7) Multi GG | 8.0 (+3.0 -2.5) | 2.0 | 2.2 (+0.9 -0.7) | 0.7 (+0.7 -0.4) | 2.2 | -0.3 |
@@
Note. — (1) Default: GG 1–10AU, 0.5–20MJup; (2) Close GG: 0.3–3AU, 0.5–20MJup; (3) Distant GG: 3–10AU, 0.5–20MJup; (4) Dynamically Hot: default GG with eccentricities 0.2; (5) Dynamically Cold: default GG with eccentricities 0.2; (6) Single GG: systems with a single gas giant in the default GG range; (7) Multi GG: systems with multiple gas giants in the default GG range. Median and 68 confidence intervals are quoted.